Could someone please help with the following question. Equations are very confusing for me.
x squared + 6x + 8 = 0
Thanks for the help.

Answer by nerdybill(7384)

The equation:
x^2 + 6x + 8 = 0
is a quadratic
.
To solve, you can either "factor" or apply the "quadratic equation".
.
Always factor if you can, it is easier.
.
x^2 + 6x + 8 = 0
(x+4)(x+2) = 0
.
This would be true if either (x+4) or (x+2) was equal to zero.
.
To find the solutions, set each to zero and solve:
x+4=0
x = -4
and
x+2=0
x = -2
.
Therefore, two solutions for x:
x = {-4,-2}